### v2

# Quantum entanglement transfer assisted via Duffing nonlinearity

Abstract
We propose a scheme to enhance quantum entanglement in an optomechanical system by exploiting the so-called Duffing nonlinearity. Our model system consists of two mechanically coupled mechanical resonators, both driven by an optical field. One resonator supports Duffing nonlinearity, while the other does not. The resonators are coupled to each other via the so-called phonon hopping mechanism. The hopping rate is $\theta$-phase-dependent that induces Exceptional Points (EPs) singularities in the system. Interestingly, while the resonator with Duffing nonlinearity exhibits vanishing bipartite entanglement, we observe an entanglement transfer phenomenon into the other mechanical resonator. This nonlinearly induced entanglement demonstrates superior robustness against thermal fluctuations compared to entanglement generated without the nonlinearity. Additionally, this entanglement features the sudden death and revival phenomenon, where the peaks happen at the multiple of $\theta=\frac{\pi}{2}$. This work opens a new avenue for exploiting nonlinear resources to generate robust quantum entanglement, paving the way for advancements in quantum information processing, quantum sensing, and quantum computing within complex systems.
Optomechanics, entanglement, exceptional point, Duffing nonlinearity, synthetic magnetism
pacs:
42.50.Wk, 42.50.Lc, 05.45.Xt, 05.45.Gg
March 22, 2024
I Introduction
Optomechanics, a field exploring the interplay between light and mechanical motion, has emerged as a captivating platform for investigating diverse physical phenomena across classical Pokharel et al. (2022); Golokolenov et al. (2021); Foulla et al. (2017); Walter and Marquardt (2016); Djorwé et al. (2013) and quantum regimes Lemonde et al. (2016); Barzanjeh et al. (2021); Aspelmeyer et al. (2010); Djorwé et al. (2012). Within the classical domain, researchers have observed numerous intriguing behaviors, including synchronization Colombano et al. (2019); Rodrigues et al. (2021); Djorwé et al. (2020); Li et al. (2022), nonlinear dynamics Djorwe et al. (2019); Navarro-Urrios et al. (2017); Roque et al. (2020); Djorwe et al. (2022), and even chaos-like states Zhu et al. (2023); Stella et al. (2023). On the other hand, the quantum realm unveils fascinating properties such as quantum correlations Purdy et al. (2017); Bemani et al. (2019); Riedinger et al. (2016) and nonclassical states Zivari et al. (2022); Fiaschi et al. (2021); Djorwé et al. (2013). These quantum effects, particularly squeezed Banerjee et al. (2023) and entangled states Riedinger et al. (2018); Kotler et al. (2021); Mercier de Lépinay et al. (2021), constitute cornerstones of quantum physics. Notably, quantum entanglement holds immense significance for various quantum technologies, including quantum information processing Meher and Sivakumar (2022); Slussarenko and Pryde (2019), quantum computing Arute et al. (2019), and high-precision sensing Xia et al. (2023).
Given the pivotal role of entanglement in numerous quantum technology applications, there is a pressing need to develop novel mechanisms for generating robust entangled states, ones that can withstand decoherence and thermal fluctuations. In this regard, researchers have exploited various nonlinear phenomena, such as the cross-Kerr effect and parametric amplification, to enhance quantum entanglement in optomechanical systems. More recently, exceptional points (EPs), non-Hermitian degeneracies, have been proposed as a tool to engineer stable and robust entanglement Chakraborty and Sarma (2019). However, synthesizing EPs in physical systems presents a significant challenge. It often necessitates precise tuning of system parameters to achieve the so-called balanced gain-loss condition, which can be experimentally demanding. This challenge has been recently addressed by the concept of synthetic magnetism, enabling the engineering of EPs even in lossy systems. This approach has paved the way for generating noise-tolerant optomechanical entanglement via dark-mode breaking Lai et al. (2022).
This work builds upon these advancements by proposing a novel scheme that combines the effects of nonlinearity and synthetic magnetism to achieve robust quantum entanglement. Our model system comprises an electromagnetic field driving two mechanically coupled resonators with a phase-dependent phonon-hopping mechanism. This phase dependence resembles synthetic magnetism, leading to the emergence of EPs singularities within the system. To introduce nonlinearity, we consider a scenario where one of the resonators exhibits Duffing nonlinearity Venkatachalam and Zhou (2023). Interestingly, our investigation reveals that while bipartite entanglement vanishes in the Duffing-nonlinear resonator, a process of entanglement transfer occurs, assisted by this nonlinearity, to the other mechanical resonator. Furthermore, the nonlinearity enhances the stability of the system, allowing for the generation of even stronger entanglement under high driving strengths. Additionally, the nonlinearly induced entanglement demonstrates superior resilience against thermal fluctuations compared to entanglement generated without the nonlinear effect. This work showcases the potential of Duffing nonlinearity as a valuable tool for enhancing quantum entanglement.
By exploring the utilization of nonlinear effects for entanglement generation, our proposal paves the way for advancements in information processing and quantum computing within complex systems. The remaining sections of this paper are structured as follows. Section II details our model, presents the associated dynamical equations, and analyzes the linear stability of the system. Section III delves into the entanglement properties and the influence of Duffing nonlinearity. Finally, Section IV concludes the paper.
II MODEL AND DYNAMICS
We consider two coupled mechanical resonators, both driven by an optical field (shown in Fig.1). The resonators are coupled through a phonon hopping $J_{m}$ mechanism, exhibiting a synthetic magnetism through the phase modulation $\theta$. The first resonator supports a Duffing nonlinearity captured by its amplitude $\eta$.
Figure 1: Schematic of the optomechanical system. An electromagnetic field (mode $a$) is driving two mechanically coupled mechanical resonators ($b_{j}$). The phonon hopping ($J_{m}$) between the two mechanical resonators is modulated through the phase $\theta$ that induces a synthetic magnetism. The optomechanical coupling between the electromagnetic field and each mechanical resonator is denoted by $g$.
Within a reference frame rotating at the driving frequency ($\omega_{p}$), the system’s Hamiltonian is described by (setting $\hbar=1$):
$\begin{split}H&=-\Delta a^{\dagger}a+\sum_{j=1,2}\omega_{j}b_{j}^{\dagger}b_{j}-ga^{\dagger}a(b_{j}+b_{j}^{\dagger})\\ &+{J_{m}}({e^{i\theta}}{b_{1}^{\dagger}}{b_{2}}+{e^{-i\theta}}{b_{1}}{b_{2}^{\dagger}})+iE^{in}(a^{\dagger}+a)\\ &+\frac{\eta}{2}(b_{1}+b_{1}^{\dagger})^{4},\end{split}$ (1)
where $a$($a^{\dagger}$) and $b_{j}$($b_{j}^{\dagger}$) represent the annihilation (creation) bosonic operators of the cavity field and the $j^{th}$ mechanical resonator (having frequency $\omega_{j}$), respectively. The detuning is defined as $\Delta=\omega_{p}-\omega_{c}$, where, $\omega_{c}$ is the cavity frequency. The other parameters are the optomechanical coupling $g$, for each resonator, and the amplitude of the driving field is $E^{in}$.
[…]
$\begin{cases}\dot{a}&=\left(\begin{array}[]{lll}i\left(\begin{array}[]{rcr}\Delta+\sum_{j=1,2}g(b_{j}^{\dagger}+b_{j})\end{array}\right)-{\frac{\kappa}{2}}\end{array}\right)a\\ &+\sqrt{\kappa}\alpha^{in}+\sqrt{\kappa}a^{in},\\ \dot{b_{1}}&=-({\frac{\gamma_{1}}{2}}+i{\omega_{1}})b_{1}-i{J_{m}}e^{i\theta}b_{2}+i{g}a^{\dagger}{a}\\ &-2i\eta(b_{1}+b_{1}^{\dagger})^{3}+\sqrt{\gamma_{1}}b_{1}^{in},\\ \dot{b_{2}}&=-({\frac{\gamma_{2}}{2}}+i{\omega_{2}})b_{2}-i{J_{m}}e^{-i\theta}b_{1}+i{g}a^{\dagger}{a}\\ &+\sqrt{\gamma_{2}}b_{2}^{in}.\end{cases}$ (2)
Here, $E^{in}$ has been substituted by $\sqrt{\kappa}\alpha^{in}$ with $\alpha^{in}$ being related to the input power $P^{in}$ as $\alpha^{in}=\sqrt{\frac{P^{in}}{\hbar\omega_{p}}}$. The cavity and mechanical dissipations are captured by $\kappa$ and $\gamma_{j}$, respectively. Moreover, $a^{in}$ and $b_{j}^{in}$ denote the zero-mean noise operators characterized by their autocorrelations functions,
$\langle a^{in}(t)a^{in\dagger}(t^{\prime})\rangle$ $=\delta(t-t^{\prime}),$ (3)
$\langle a^{in\dagger}(t)a^{in}(t^{\prime})\rangle$ $=0,$ (4)
$\langle b_{j}^{in}(t)b_{j}^{in\dagger}(t^{\prime})\rangle$ $=(n_{th}^{j}+1)\delta(t-t^{\prime}),$ (5)
$\langle b_{j}^{in\dagger}(t)b_{j}^{in}(t^{\prime})\rangle$ $=n_{th}^{j}\delta(t-t^{\prime}),$ (6)
with $n_{th}^{j}=\left[\rm exp\left(\frac{\hbar\omega_{j}}{k_{b}T}\right)-1\right]^{-1}$ representing the number of phonons at temperature $T$ and $k_{B}$ the Boltzmann constant.
To obtain the system’s quadratures, the nonlinear set of equations (2) can be linearized through a standard linearization process. This involves splitting operators into their mean values and fluctuations, $\mathcal{O}=\langle\mathcal{O}\rangle+\delta\mathcal{O}$, where $\mathcal{O}\equiv$ ($a$,$b_{j}$). By setting $\alpha=\langle a\rangle$, and $\beta_{j}=\langle b_{j}\rangle$), the dynamical mean values of our system yields,
$\begin{cases}\dot{\alpha}&=\left(i\tilde{\Delta}-\frac{\kappa}{2}\right)\alpha+\sqrt{\kappa}\alpha^{in},\\ \dot{\beta}_{1}&=-({\frac{\gamma_{1}}{2}}+i{\omega_{1}}){\beta}_{1}-i{J_{m}}e^{i\theta}{\beta}_{2}+i{g}|\alpha|^{2}\\ &-2i\eta(\beta_{1}+\beta_{1}^{*})^{3},\\ \dot{\beta_{2}}&=-({\frac{\gamma_{2}}{2}}+i{\omega_{2}}){\beta}_{2}-i{J_{m}}e^{-i\theta}{\beta}_{1}+i{g}|\alpha|^{2},\end{cases}$ (7)
and the fluctuation dynamics reads,
$\begin{cases}\delta\dot{a}&=\left(i\tilde{\Delta}-\frac{\kappa}{2}\right)\delta a+i\sum_{j=1,2}G({\delta}b_{j}^{\dagger}+{\delta}b_{j})\\ &+\sqrt{\kappa}a^{in},\\ \delta\dot{b_{1}}&=-({\frac{\gamma_{1}}{2}}+i{\omega_{1}})\delta{b}_{1}-i{J_{m}}e^{i\theta}\delta{b}_{2}+i(G^{*}\delta a\\ &+G\delta a^{\dagger})-i\Lambda(\delta b_{1}+\delta b_{1}^{\dagger})+\sqrt{\gamma_{1}}b_{1}^{in},\\ \delta\dot{b_{2}}&=-({\frac{\gamma_{2}}{2}}+i{\omega_{2}})\delta{b}_{2}-i{J_{m}}e^{-i\theta}\delta{b}_{1}+i(G^{*}\delta a\\ &+G\delta a^{\dagger})+\sqrt{\gamma_{2}}b_{2}^{in},\end{cases}$ (8)
where both the effective detuning and the optomechanical coupling are defined respectively as: $\tilde{\Delta}=\Delta+2\sum_{j=1,2}g\rm Re(\beta_{j})$ and $G=g\alpha$. From now on, we will assume that $\alpha$ is real, meaning that $G$ is real as well. Furthermore, $\Lambda=24\eta(\Re(\beta_{1}))^{2}$ captures the nonlinear process. The fluctuation dynamics in Eq.(8) can be put in its compact form,
$\dot{x}=\rm{A}x+y,$ (9)
where ${x}=(\delta a,\delta a^{\dagger},\delta b_{1},\delta b^{\dagger}_{1},\delta b_{2},\delta b^{\dagger}_{2})^{T}$, and the noise vector $y=(\sqrt{\kappa}a^{in},\sqrt{\kappa}a^{in\dagger},\sqrt{\gamma_{1}}b_{1}^{in},\sqrt{\gamma_{1}}b^{in\dagger}_{1},\sqrt{\gamma_{2}}b_{2}^{in},\sqrt{\gamma_{2}}b^{in\dagger}_{2})^{T}$. The matrix $\rm{A}$ is given by:
$\rm{A}=\begin{pmatrix}A_{1}&A_{12}&A_{13}\\ A_{12}^{T}&A_{2}&A_{23}\\ A_{13}^{T}&A_{23}^{T}&A_{3}\end{pmatrix},$ (10)
with
$A_{1}$ $=\begin{pmatrix}i\tilde{\Delta}-\frac{\kappa}{2}&0\\ 0&-i\tilde{\Delta}-\frac{\kappa}{2}\end{pmatrix},$ (13)
$A_{2}$ $=\begin{pmatrix}-(\frac{\gamma_{1}}{2}+i(\omega_{1}+\Lambda))&-i\Lambda\\ i\Lambda&-(\frac{\gamma_{1}}{2}-i(\omega_{1}+\Lambda))\end{pmatrix},$ (16)
$A_{3}$ $=\begin{pmatrix}m-(\frac{\gamma_{2}}{2}+i\omega_{2})&0\\ 0&-(\frac{\gamma_{2}}{2}-i\omega_{2})\end{pmatrix},$ (19)
$A_{12}$ $=\begin{pmatrix}iG&iG\\ -iG^{*}&-iG^{*}\end{pmatrix},$ (22)
$A_{13}$ $=\begin{pmatrix}iG&iG\\ -iG^{*}&-iG^{*}\end{pmatrix},$ (25)
$A_{23}$ $=\begin{pmatrix}-iJ_{m}e^{i\theta}&0\\ 0&iJ_{m}e^{-i\theta}\end{pmatrix}.$ (28)
The system’s steady-state solution is achieved when the mean values in Eq.(7) become time-independent ($\dot{\alpha}=\dot{\beta_{j}}=0$). Stability diagrams, shown in Fig. 2, are then generated using the Routh-Hurwitz criterion and the steady-state values within the A matrix.
Figure 2: Stability diagrams. (a) Driving strength $\alpha^{in}$ vs the non-linear amplitude $\eta$ for $\kappa=2\times 10^{-1}$ and $J_{m}=1\times 10^{-2}$. (b) Driving strength $\alpha^{in}$ vs the cavity decay rate $\kappa$ for $\eta=1\times 10^{-5}$ and $J_{m}=1\times 10^{-2}$. The dark color is stable, while the light color is unstable. The other parameters are $\omega_{1}=\omega_{2}=\omega_{m}$, $\gamma_{1}=\gamma_{2}=1\times 10^{-5}\omega_{m}$, $\Delta=-\omega_{m}$, $g=5\times 10^{-4}\omega_{m}$ and $\theta=\frac{\pi}{2}$.
Fig.2(a) shows the stability diagram in ($\eta,\alpha^{in}$) space. It could be seen that for a wide range of the Duffing nonlinear parameters, the system is stable for a driving strength up to $\alpha^{in}\sim 1500\omega_{m}^{1/2}$. It can also be seen that the stability of the system improves slightly as the nonlinearity increases. For $\eta=10^{-5}\omega_{m}$, Fig.2(b) displays stability of the system in ($\kappa,\alpha^{in}$) space. It can clearly be seen that the system is stable for a wide range of $\kappa$. Here we have considered the synthetic phase to be, $\theta\sim\pi/2$, which corresponds to the optimal value of $\theta$ around which the logarithmic negativity is strong enough, as could be seen later. For our study of quantum entanglement, we assume that the given system fulfills the stability conditions depicted in Fig.2.
III Stationary entanglement transfer
We define the quantum quadratures as follows:
$\delta{{x}}=\frac{\delta{a}^{\dagger}+\delta{a}}{\sqrt{2}}$, $\delta{{y}}=i\frac{\delta{a}^{\dagger}-\delta{a}}{\sqrt{2}}$, $\delta{q}_{j}=\frac{\delta{b}_{j}^{\dagger}+\delta{b}_{j}}{\sqrt{2}}$, $\delta{{p}}_{j}=i\frac{\delta{{b}}_{j}^{\dagger}-\delta{b}_{j}}{\sqrt{2}}$,
with their corresponding noise operators $\delta{{x}}^{in}=\frac{\delta{a}^{in\dagger}+\delta{a}^{in}}{\sqrt{2}}$, $\delta{{y}}^{in}=i\frac{\delta{a}^{in\dagger}-\delta{a}^{in}}{\sqrt{2}}$, $\delta{q}_{j}^{in}=\frac{\delta{b}_{j}^{in\dagger}+\delta{b}_{j}^{in}}{\sqrt{2}}$ and $\delta{{p}}_{j}^{in}=i\frac{\delta{{b}}_{j}^{in\dagger}-\delta{b}_{j}^{in}}{\sqrt{2}}$. By using Eq.(8), the system quadratures can be written in its compact form as,
$\dot{u}=\rm{M}u+z,$ (29)
where the quadratures vector is ${u}^{T}=(\delta{x},\delta{y},\delta{q}_{1},\delta{p}_{1},\delta{q}_{2},\delta{p}_{2})$, with the corresponding noise vector set as $z^{T}=(\sqrt{\kappa}x^{in},\sqrt{\kappa}y^{in},\sqrt{\gamma_{1}}q_{1}^{in},\sqrt{\gamma_{1}}p_{1}^{in},\sqrt{\gamma_{2}}q_{2}^{in},\sqrt{\gamma_{2}}p_{2}^{in})$. The drift matrix is:
$\rm{M}=\begin{pmatrix}M_{1}&M_{12}&M_{13}\\ M_{12}^{T}&M_{2}&M_{23}\\ M_{13}^{T}&M_{23}^{T}&M_{3}\end{pmatrix},$ (30)
where $M_{i}$ and $M_{ij}$ are blocs of $2\times 2$ matrices defined as,
$M_{1}$ $=\begin{pmatrix}-\frac{\kappa}{2}&-\tilde{\Delta}\\ \tilde{\Delta}&-\frac{\kappa}{2}\end{pmatrix},$ (33)
$M_{2}$ $=\begin{pmatrix}-\frac{\gamma_{1}}{2}&\omega_{1}\\ -(2\Lambda+\omega_{1})&-\frac{\gamma_{1}}{2}\end{pmatrix},$ (36)
$M_{3}$ $=\begin{pmatrix}-\frac{\gamma_{2}}{2}&\omega_{2}\\ -\omega_{2}&-\frac{\gamma_{2}}{2}\end{pmatrix},$ (39)
$M_{12}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ -2\rm Re(G)&0\end{pmatrix},$ (42)
$M_{13}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ -2\rm Re(G)&0\end{pmatrix},$ (45)
$M_{23}$ $=\begin{pmatrix}J_{m}\sin\theta&J_{m}\cos\theta\\ -J_{m}\cos\theta&J_{m}\sin\theta\end{pmatrix}.$ (48)
Figure 3: (a) Entanglement between the electromagnetic field and the nonlinear resonator (mode $b_{1}$) versus the synthetic phase $\theta$. (b) Entanglement between the electromagnetic field and the linear resonator (mode $b_{2}$) versus the synthetic phase $\theta$. The solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively. We have used $\kappa=2\times 10^{-1}\omega_{m}$, $J_{m}=2\times 10^{-1}\omega_{m}$, $\alpha^{in}=1500$ and $n_{th}=100$. The other parameters are as in Fig.2.
The diagonal blocks denoted as $M_{i}$ for $i=1,2,3$ correspond to the optical mode ($i=1$), the first mechanical mode ($i=2$), and the second mechanical mode ($i=3$), respectively. The off-diagonal blocks capture the correlations between different system components: $M_{12}$ and $M_{13}$ describe the correlations between the driving field and the mechanical resonators, while $M_{23}$ represents the correlations between the two mechanical modes themselves. Bipartite entanglement within this system can be quantified using the logarithmic negativity ($\rm{E_{N}}$), which can be evaluated by tracing out the non-necessary third mode. $\rm{E_{N}}$ is defined as,
$E_{N}=\rm max[0,-\ln(2\nu^{-})],$ (49)
where $\nu^{-}=\frac{1}{\sqrt{2}}\sqrt{\sum(V)-\sqrt{\sum(V)^{2}-4\rm detV}}$. The covariance matrix $V$ contains elements defined as,
$V_{ij}=\frac{\left(\langle u_{i}u_{j}+u_{j}u_{i}\rangle\right)}{2},$ (50)
where $u_{i}$ and $u_{j}$ represent quadratures of the system’s operators. Notably, $V$ can be expressed as a standard $2\times 2$ matrix,
$\rm{V}=\begin{pmatrix}A&C\\ C^{T}&B\end{pmatrix},$ (51)
such that $\sum(V)=\rm{detA+detB-2detC}$.
Under the condition of system stability, the covariance matrix $V$ satisfies the Lyapunov equation,
$MV+VM^{T}=-D,$ (52)
where $D=\rm{Diag}[\frac{\kappa}{2},\frac{\kappa}{2},\frac{\gamma_{1}}{2}(2n_{th}+1),\frac{\gamma_{1}}{2}(2n_{th}+1),\frac{\gamma_{2}}{2}(2n_{th}+1),\frac{\gamma_{2}}{2}(2n_{th}+1)]$. The diagonal elements of $D$, $D_{ii}$, represent the individual noise terms for each system component. These terms include the cavity decay rate ($\kappa$) for the optical mode and the mechanical dissipation rates ($\gamma_{i}$) for the two mechanical resonators. Notably, each term is scaled by a factor of ($12n_{th}+$ 1). This factor accounts for the contribution of thermal noise to the overall dissipation within the system. The matrix $D$ is referred to as the matrix of noise correlations. This terminology emphasizes that the diagonal elements $D_{ii}$ are not simply independent noise sources, but rather they represent the inherent correlations between these noise terms due to the system’s thermal environment.
Fig.3 captures how the bipartite entanglement between the driving field and each mechanical resonator varies with the synthetic phase $\theta$. The interaction with the first mechanical resonator is depicted in Fig.3(a), while entanglement involving the second mechanical resonator is shown in Fig.3(b). In these figures, the solid line represents the case without the Duffing nonlinearity ($\eta=0$), while the dash-dotted and the dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively. For $\eta=0$, these figures show how the entanglement is modulated by the synthetic phase. Indeed, the entanglement $\rm{E_{N}}$ drops around $\theta=n\pi$ for $n\in\mathbb{N}$, and it reaches its optimal value at the vicinity of $\theta=m\pi/2$ where $m$ is and odd number. This dynamical behavior of the entanglement is reminiscent of sudden death and revival of entanglement induced through exceptional points in Chakraborty and Sarma (2019). Interestingly, for our system with $\eta=0$, the synthetic magnetism phase might also harbor exceptional points for $\theta=m\pi/2$. When $\eta\neq 0$, Fig. 3(a) suggests a general weakening of the entanglement between the driving field and the first mechanical resonator (supporting the Duffing nonlinearity). Conversely, Fig. 3(b) indicates an enhancement of entanglement with the second mechanical resonator (free of the Duffing term). This implies that the Duffing nonlinearity tends to suppress entanglement on the resonator it resides in, while transferring it towards the connected, nonlinearity-free resonator.
Figure 4: Entanglement between the optical field with the nonlinear resonator (a), and with the linear resonator (b) versus the driving field $\alpha^{in}$. Entanglement between the optical field with the nonlinear resonator (c), and with the linear resonator (d) versus the cavity decay rate $\kappa$. In all these figures, the solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively. The phonon number is $n_{th}^{1,2}=100$ for all figures, while in (c,d) $\alpha^{in}=1000$. The rest of the parameters are as in Fig.3.
Figures 4 delve deeper into the influence of the Duffing nonlinearity by examining how entanglement between the optical field and each mechanical resonator varies with the driving field strength $\alpha^{in}$ (Figs.4 (a,b)) and the cavity decay rate $\kappa$ (Figs.4(c,d)). The solid line again represents the case without the nonlinearity ($\eta=0$). As the nonlinear parameter ($\eta$) increases, the key observation is a decrease and eventual vanishing of entanglement with the nonlinear resonator (Figs. 4(a,c)). Conversely, there is a corresponding enhancement of entanglement with the nonlinearity-free resonator (Figs. 4(b,d)). Beyond this entanglement transfer induced by the Duffing nonlinearity, Fig. 2(a) also suggests that this nonlinearity improves system stability. This stability improvement, in turn, allows for entanglement to persist over a wider range of parameter values compared to the case where the resonators lack the Duffing term. Therefore, these results highlight the potential of nonlinear effects for manipulating and enhancing entanglement in coupled resonators.
Figure 5: Entanglement between the optical field with the nonlinear resonator (a), and with the linear resonator (b) versus the nonlinear term $\eta$. Entanglement between the electromagnetic field with the nonlinear resonator (c), and with the linear resonator (d) versus the thermal phonon number $n_{th}$. In (a,b), the solid line is for $n_{th=0}$ and the dashed line is for $n_{th=100}$. In (c,d), the solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively. The driving field is $\alpha^{in}=1500$ in (a,b) and $\alpha^{in}=1000$ in (c,d). The rest of the parameters are as in Fig.3.
Figures 5 investigate the robustness and stability of entanglement against thermal fluctuations. Figure 5(a) depicts the entanglement behavior with respect to the Duffing nonlinearity for the resonator supporting the nonlinearity, while Fig. 5(b) shows the same information for the nonlinearity-free resonator. The solid line in these figures represents the case with no thermal noise ($n_{th}=0$), while the dashed line corresponds to a finite thermal phonon number ($n_{th}=100$). Examining the solid lines confirms that the Duffing nonlinearity enhances entanglement. While some decrease in entanglement strength is observed for the nonlinearity-free resonator in the presence of thermal fluctuations, the transferred entanglement exhibits a superior level of robustness compared to that generated directly on the Duffing resonator (compare the solid lines in Fig. 5(a) and (b)). The robustness of the transferred entanglement is further emphasized in Figs. 5(c,d), which depict the logarithmic negativity as a function of the thermal phonon number. By comparing the dashed and dash-dotted lines in these figures, it becomes clear that the transferred entanglement exhibits greater resilience against thermal fluctuations as the nonlinearity strength increases. These findings demonstrate how Duffing nonlinearity can be harnessed to promote strong and thermally stable entanglement, leading to potential advancements in quantum information processing and quantum computing within complex systems that incorporate nonlinearities.
IV CONCLUSION
In this work, we investigated stationary entanglement in an optomechanical system involving a single optical field driving two coupled mechanical resonators. The inter-resonator phonon hopping rate was modulated to induce a synthetic magnetism effect. Additionally, one of the resonators incorporated a Duffing nonlinearity. Utilizing experimentally achievable parameters, we demonstrated a slight improvement in system stability, a favorable condition for enhancing stationary entanglement. For the case without the Duffing nonlinearity, the logarithmic negativity exhibited the phenomenon of sudden death and revival of entanglement. The entanglement reached optimal values near the synthetic phase of $\theta=m\frac{\pi}{2}$ (where $m$ is an odd integer) and vanished around $\theta=n\pi$ (where $n$ is any integer). Following a strategic selection of the synthetic phase, we observed a weakening and eventual vanishing of entanglement between the optical field and the Duffing resonator when the nonlinearity became sufficiently strong. However, there was a corresponding enhancement of entanglement with the nonlinearity-free resonator. This behavior suggests a process of entanglement transfer facilitated by the Duffing nonlinearity. Notably, the transferred entanglement displayed superior robustness against thermal fluctuations. Our findings pave the way for further exploration of entanglement transfer phenomena in complex systems. These results hold promise for generating robust and stable entanglement in nonlinear systems, which could prove instrumental for advancements in quantum information processing and quantum computing applications.
Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. A.K. Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161).

### v3

@@
-Quantum entanglement transfer assisted via Duffing nonlinearity
+Quantum entanglement assisted via Duffing nonlinearity
 Abstract
@@
 Our model system consists of two mechanically coupled mechanical resonators, both driven by an optical field.
-One resonator supports Duffing nonlinearity, while the other does not.
-The resonators are coupled to each other via the so-called phonon hopping mechanism.
+One resonator supports Duffing nonlinearity, while the other does not.The resonators are coupled to each other via the so-called phonon hopping mechanism.
 The hopping rate is $\theta$-phase-dependent that induces Exceptional Points (EPs) singularities in the system.
-Interestingly, while the resonator with Duffing nonlinearity exhibits vanishing bipartite entanglement, we observe an entanglement transfer phenomenon into the other mechanical resonator.
-This nonlinearly induced entanglement demonstrates superior robustness against thermal fluctuations compared to entanglement generated without the nonlinearity.
-Additionally, this entanglement features the sudden death and revival phenomenon, where the peaks happen at the multiple of $\theta=\frac{\pi}{2}$.
-This work opens a new avenue for exploiting nonlinear resources to generate robust quantum entanglement, paving the way for advancements in quantum information processing, quantum sensing, and quantum computing within complex systems.
+Interestingly, while the resonator with Duffing nonlinearity exhibits vanishing entanglement with light, we observe an increase in entanglement between light and the other mechanical resonator.
+This enhanced entanglement persists longer against thermal fluctuations compared to the one without the nonlinearity.
+Additionally, this entanglement features a sudden death and revival phenomenon, where the peaks happen at the multiple of $\theta=\frac{\pi}{2}$.
+This work opens a new avenue for exploiting nonlinear resources to generate strong quantum entanglement, paving the way for advancements in quantum information processing, quantum sensing, and quantum computing within complex systems.
 Optomechanics, entanglement, exceptional point, Duffing nonlinearity, synthetic magnetism
@@
 42.50.Wk, 42.50.Lc, 05.45.Xt, 05.45.Gg
-March 22, 2024
+October 9, 2024
 I Introduction
@@
 (2013).
-These quantum effects, particularly squeezed Banerjee et al.
+These quantum effects, particularly squeezed Lü et al.
+(2015); Banerjee et al.
 (2023) and entangled states Riedinger et al.
+(2018); Chen et al.
 (2020); Kotler et al.
@@
 Given the pivotal role of entanglement in numerous quantum technology applications, there is a pressing need to develop novel mechanisms for generating robust entangled states, ones that can withstand decoherence and thermal fluctuations.
-In this regard, researchers have exploited various nonlinear phenomena, such as the cross-Kerr effect and parametric amplification, to enhance quantum entanglement in optomechanical systems.
-More recently, exceptional points (EPs), non-Hermitian degeneracies, have been proposed as a tool to engineer stable and robust entanglement Chakraborty and Sarma (2019).
+In this regard, researchers have exploited various nonlinear phenomena, such as the cross-Kerr effect, parametric amplification, and Duffing nonlinearity Djorwe et al.
+(2014) to enhance quantum entanglement in optomechanical systems.
+These nonlinear effect enhancing entanglement where realized in single optomechanical cavities.
+Owing to the advance in nanofabrication, the engineering of coupled optomechanical cavities has been possible.
+The aim behind such scalability of optomechanical structures is the improvement of the optomechanical coupling, which is a requirement of strong entanglement.
+More recently, such coupled optomechanical structures have led to exceptional points (EPs), non-Hermitian degeneracies, which have been proposed as a tool to engineer stable and robust entanglement Chakraborty and Sarma (2019).
 However, synthesizing EPs in physical systems presents a significant challenge.
@@
 (2022).
-This work builds upon these advancements by proposing a novel scheme that combines the effects of nonlinearity and synthetic magnetism to achieve robust quantum entanglement.
+Moreover, these EPs are good tools to engineer the sudden death and revival of entanglement phenomenon, which could be used to tune quantum correlations.
+This work builds upon these advancements by proposing a novel scheme that combines the effects of nonlinearity and synthetic magnetism to achieve robust quantum entanglement between electromagnetic field and mechanical resonators.
 Our model system comprises an electromagnetic field driving two mechanically coupled resonators with a phase-dependent phonon-hopping mechanism.
@@
 To introduce nonlinearity, we consider a scenario where one of the resonators exhibits Duffing nonlinearity Venkatachalam and Zhou (2023).
-Interestingly, our investigation reveals that while bipartite entanglement vanishes in the Duffing-nonlinear resonator, a process of entanglement transfer occurs, assisted by this nonlinearity, to the other mechanical resonator.
+Interestingly, our investigation reveals that while entanglement vanishes in the Duffing-nonlinear resonator, there is an entanglement induced to the other mechanical resonator that is assisted by this nonlinearity.
 Furthermore, the nonlinearity enhances the stability of the system, allowing for the generation of even stronger entanglement under high driving strengths.
@@
 The resonators are coupled through a phonon hopping $J_{m}$ mechanism, exhibiting a synthetic magnetism through the phase modulation $\theta$.
+The phase $\theta$ can be also tuned by adjusting the phase of the electromagnetic fields acting on the mechanical resonators involved as shown in Chen et al.
+(2023); Lai et al.
+(2022).
 The first resonator supports a Duffing nonlinearity captured by its amplitude $\eta$.
-Figure 1: Schematic of the optomechanical system.
+It is worthwhile to mention that such a system, driven by a microwave, could be realized experimentally in circuit electromechanical system (refer to Fig.1(a)) .
+Here, the phonon-hopping mechanism could be engineered between the nearest-neighboring micromechanical resonator using a superconducting circuit Lai et al.
+(2022).
+In Fig.1(b), we depict a possible optomechanical version of the system.
+Figure 1: Sketch of our optomechanical system.
+(a) A microwave optomechanical version of our system, where a microwave optical field is generated by an LC resonator and is driving two coupled mechanical resonators.
+(b) A possible optomechanical version of our system, where an optical field is driving two mechanically coupled mechanical resonators.
+The mechanical displacement are $x_{1}$ and $x_{2}$, respectively.
+(c) A three-mode representation of our system.
 An electromagnetic field (mode $a$) is driving two mechanically coupled mechanical resonators ($b_{j}$).
 The phonon hopping ($J_{m}$) between the two mechanical resonators is modulated through the phase $\theta$ that induces a synthetic magnetism.
+The first mode supports a Duffing nonlinearity $\eta$.
 The optomechanical coupling between the electromagnetic field and each mechanical resonator is denoted by $g$.
-Within a reference frame rotating at the driving frequency ($\omega_{p}$), the system’s Hamiltonian is described by (setting $\hbar=1$):
+Within a reference frame rotating at the driving frequency ($\omega_{p}$), the system’s Hamiltonian is described by (setting $\hbar=1$) Lai et al. (2022):
 $\begin{split}H&=-\Delta a^{\dagger}a+\sum_{j=1,2}\omega_{j}b_{j}^{\dagger}b_{j}-ga^{\dagger}a(b_{j}+b_{j}^{\dagger})\\ &+{J_{m}}({e^{i\theta}}{b_{1}^{\dagger}}{b_{2}}+{e^{-i\theta}}{b_{1}}{b_{2}^{\dagger}})+iE^{in}(a^{\dagger}+a)\\ &+\frac{\eta}{2}(b_{1}+b_{1}^{\dagger})^{4},\end{split}$ (1)
 where $a$($a^{\dagger}$) and $b_{j}$($b_{j}^{\dagger}$) represent the annihilation (creation) bosonic operators of the cavity field and the $j^{th}$ mechanical resonator (having frequency $\omega_{j}$), respectively.
+The operators $b_{j}$ are related to the displacements $x_{j}$ through $x_{1}=X_{ZPF}(b_{j}+b_{j}^{\dagger})$, where $X_{ZPF}$ is the mechanical zero point motion of the mechanical resonators.
 The detuning is defined as $\Delta=\omega_{p}-\omega_{c}$, where, $\omega_{c}$ is the cavity frequency.
@@
 Stability diagrams, shown in Fig.
-2, are then generated using the Routh-Hurwitz criterion and the steady-state values within the A matrix.
+2, are then generated using the Routh-Hurwitz criterion DeJesus and Kaufman (1987) and the steady-state values within the A matrix.
 Figure 2: Stability diagrams.
-(a) Driving strength $\alpha^{in}$ vs the non-linear amplitude $\eta$ for $\kappa=2\times 10^{-1}$ and $J_{m}=1\times 10^{-2}$.
-(b) Driving strength $\alpha^{in}$ vs the cavity decay rate $\kappa$ for $\eta=1\times 10^{-5}$ and $J_{m}=1\times 10^{-2}$.
-The dark color is stable, while the light color is unstable.
+(a) Driving strength $\alpha^{in}$ vs the non-linear amplitude $\eta$ for $\kappa=2\times 10^{-1}\omega_{m}$ and $J_{m}=1\times 10^{-2}\omega_{m}$.
+(b) Driving strength $\alpha^{in}$ vs the cavity decay rate $\kappa$ for $\eta=1\times 10^{-5}\omega_{m}$ and $J_{m}=1\times 10^{-2}\omega_{m}$.
+The blue color is stable, while the red color is unstable.
 The other parameters are $\omega_{1}=\omega_{2}=\omega_{m}$, $\gamma_{1}=\gamma_{2}=1\times 10^{-5}\omega_{m}$, $\Delta=-\omega_{m}$, $g=5\times 10^{-4}\omega_{m}$ and $\theta=\frac{\pi}{2}$.
+In the rest of the paper, for our analysis, we have used the following experimentally feasible parameters Fang et al.
+(2017); Mercier de Lépinay et al.
+(2021): $\omega_{1}=\omega_{2}=\omega_{m}$, $\gamma_{1}=\gamma_{2}=1\times 10^{-5}\omega_{m}$, $\Delta=-\omega_{m}$, $g=5\times 10^{-4}\omega_{m}$, $\kappa=2\times 10^{-1}\omega_{m}$ and $J_{m}=1\times 10^{-2}\omega_{m}$, and $\theta=\frac{\pi}{2}$.
 Fig.2(a) shows the stability diagram in ($\eta,\alpha^{in}$) space.
@@
 For our study of quantum entanglement, we assume that the given system fulfills the stability conditions depicted in Fig.2.
-III Stationary entanglement transfer
+III Stationary entanglement generation and Duffing nonlinear effect
 We define the quantum quadratures as follows:
@@
 $M_{3}$ $=\begin{pmatrix}-\frac{\gamma_{2}}{2}&\omega_{2}\\ -\omega_{2}&-\frac{\gamma_{2}}{2}\end{pmatrix},$ (21)
-$M_{12}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ -2\rm Re(G)&0\end{pmatrix},$ (42)
-$M_{13}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ -2\rm Re(G)&0\end{pmatrix},$ (45)
+$M_{12}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ 2\rm Re(G)&0\end{pmatrix},$ (22)
+$M_{13}$ $=\begin{pmatrix}-2\rm Im(G)&0\\ 2\rm Re(G)&0\end{pmatrix},$ (23)
 $M_{23}$ $=\begin{pmatrix}J_{m}\sin\theta&J_{m}\cos\theta\\ -J_{m}\cos\theta&J_{m}\sin\theta\end{pmatrix}.$ (24)
@@
 The solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively.
-We have used $\kappa=2\times 10^{-1}\omega_{m}$, $J_{m}=2\times 10^{-1}\omega_{m}$, $\alpha^{in}=1500$ and $n_{th}=100$.
+We have used $\kappa=2\times 10^{-1}\omega_{m}$, $J_{m}=2\times 10^{-1}\omega_{m}$, $\alpha^{in}=1000\omega_{m}^{1/2}$ and $n_{th}=100$.
 The other parameters are as in Fig.2.
@@
 Bipartite entanglement within this system can be quantified using the logarithmic negativity ($\rm{E_{N}}$), which can be evaluated by tracing out the non-necessary third mode.
-$\rm{E_{N}}$ is defined as,
-$E_{N}=\rm max[0,-\ln(2\nu^{-})],$ (49)
+$\rm{E_{N}}$ is defined as Djorwe et al.
+(2014); Lai et al.
+(2022),
+$\rm{E_{N}}=\rm max[0,-\ln(2\nu^{-})],$ (25)
 where $\nu^{-}=\frac{1}{\sqrt{2}}\sqrt{\sum(V)-\sqrt{\sum(V)^{2}-4\rm detV}}$. The covariance matrix $V$ contains elements defined as,
@@
 $MV+VM^{T}=-D,$ (28)
-where $D=\rm{Diag}[\frac{\kappa}{2},\frac{\kappa}{2},\frac{\gamma_{1}}{2}(2n_{th}+1),\frac{\gamma_{1}}{2}(2n_{th}+1),\frac{\gamma_{2}}{2}(2n_{th}+1),\frac{\gamma_{2}}{2}(2n_{th}+1)]$.
+where $D=\rm{Diag}[\frac{\kappa}{2},\frac{\kappa}{2},\frac{\gamma_{1}}{2}(2n_{th}^{1}+1),\frac{\gamma_{1}}{2}(2n_{th}^{1}+1),\frac{\gamma_{2}}{2}(2n_{th}^{2}+1),\frac{\gamma_{2}}{2}(2n_{th}^{2}+1)]$.
 The diagonal elements of $D$, $D_{ii}$, represent the individual noise terms for each system component.
@@
 3(b) indicates an enhancement of entanglement with the second mechanical resonator (free of the Duffing term).
-This implies that the Duffing nonlinearity tends to suppress entanglement on the resonator it resides in, while transferring it towards the connected, nonlinearity-free resonator.
+This implies that the Duffing nonlinearity tends to suppress entanglement on the resonator it resides in, while inducing it towards the connected, nonlinearity-free resonator.
 Figure 4: Entanglement between the optical field with the nonlinear resonator (a), and with the linear resonator (b) versus the driving field $\alpha^{in}$.
@@
 In all these figures, the solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively.
-The phonon number is $n_{th}^{1,2}=100$ for all figures, while in (c,d) $\alpha^{in}=1000$.
+The phonon number is $n_{th}^{1,2}=100$ for all figures, while in (c,d) $\alpha^{in}=1000\omega_{m}^{1/2}$.
 The rest of the parameters are as in Fig.3.
@@
 4(b,d)).
-Beyond this entanglement transfer induced by the Duffing nonlinearity, Fig.
+Beyond this entanglement induced by the Duffing nonlinearity, Fig.
 2(a) also suggests that this nonlinearity improves system stability.
@@
 Therefore, these results highlight the potential of nonlinear effects for manipulating and enhancing entanglement in coupled resonators.
+It is worthwhile to note that, in our simulations and for the range of our used parameters, there is no regime where the nonlinearity improves or reduces entanglement of the cavity mode with both the mechanical resonators.
+Also, for strong enough value of the nonlinear Duffing parameter,$\eta$, the system becomes unstable.
+On the other hand, for a negative value of the nonlinear term, our system is always unstable, and no entanglement is generated.
 Figure 5: Entanglement between the optical field with the nonlinear resonator (a), and with the linear resonator (b) versus the nonlinear term $\eta$.
@@
 In (c,d), the solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-5}\omega_{m}$, respectively.
-The driving field is $\alpha^{in}=1500$ in (a,b) and $\alpha^{in}=1000$ in (c,d).
+The driving field is $\alpha^{in}=1500\omega_{m}^{1/2}$ in (a,b) and $\alpha^{in}=1000\omega_{m}^{1/2}$ in (c,d).
 The rest of the parameters are as in Fig.3.
-Figures 5 investigate the robustness and stability of entanglement against thermal fluctuations.
+Figures 5 investigate how our generated entanglement resists against thermal fluctuations.
 Figure 5(a) depicts the entanglement behavior with respect to the Duffing nonlinearity for the resonator supporting the nonlinearity, while Fig.
@@
 Examining the solid lines confirms that the Duffing nonlinearity enhances entanglement.
-While some decrease in entanglement strength is observed for the nonlinearity-free resonator in the presence of thermal fluctuations, the transferred entanglement exhibits a superior level of robustness compared to that generated directly on the Duffing resonator (compare the solid lines in Fig.
+While some decrease in entanglement strength is observed for the nonlinearity-free resonator in the presence of thermal fluctuations, the induced entanglement exhibits a superior level of resilience compared to that generated directly on the Duffing resonator (compare the solid lines in Fig.
 5(a) and (b)).
-The robustness of the transferred entanglement is further emphasized in Figs.
+The resilience of the induced entanglement is further emphasized in Figs.
 5(c,d), which depict the logarithmic negativity as a function of the thermal phonon number.
-By comparing the dashed and dash-dotted lines in these figures, it becomes clear that the transferred entanglement exhibits greater resilience against thermal fluctuations as the nonlinearity strength increases.
+By comparing the dashed and dash-dotted lines in these figures, it becomes clear that the induced entanglement exhibits greater resilience against thermal fluctuations as the nonlinearity strength increases.
 These findings demonstrate how Duffing nonlinearity can be harnessed to promote strong and thermally stable entanglement, leading to potential advancements in quantum information processing and quantum computing within complex systems that incorporate nonlinearities.
+The physics behind why entanglement is quite robust against noise, and why Duffing nonlinearity improves entanglement could be attributed to the dark-mode breaking phenomena induced by synthetic magnetism and Duffing nonlinearity.
+In Lai et al.
+(2022), the dark-mode breaking phenomena due to synthetic magnetism is discussed at great length.
+While the same context is applicable for our proposed scheme, here we give a brief intuitive explanation.
+In the absence of synthetic magnetism and Duffing nonlinearity (i.e.
+$J_{m}=0$ and $\Lambda=0$ or $\eta=0$), it is possible to show (refer to Eq.8) that the system possess two hybrid mechanical bright ($B$) and dark ($D$) modes, defined by: $B=(\delta b_{1}+\delta b_{2})/\sqrt{2}$ and $D=(\delta b_{1}-\delta b_{2})/\sqrt{2}$.
+The dark mode, $D$, destroys all quantum resources as it is decoupled from the system.
+However, this dark-mode could be broken, i.e.
+it could be coupled to the system by introducing synthetic magnetism (i.e.
+when $J_{m}\neq 0$).
+This breaking is further enhanced by incorporating Duffing nonlinearity(i.e.$\Lambda\neq 0$).
+Also, for $\theta=n\pi$, $n$ is an integer, either $B$ or $D$ become a dark mode.
+Hence, tuning $\theta\neq n\pi$, the dark mode could be coupled to the optical mode.
+This explains the reason why we have an optimal value of entanglement around $\theta=m\pi/2$ ($m$ an odd integer), even if there is no Duffing nonlinearity (i.e.
+$\eta=0$).
 IV CONCLUSION
@@
 However, there was a corresponding enhancement of entanglement with the nonlinearity-free resonator.
-This behavior suggests a process of entanglement transfer facilitated by the Duffing nonlinearity.
-Notably, the transferred entanglement displayed superior robustness against thermal fluctuations.
-Our findings pave the way for further exploration of entanglement transfer phenomena in complex systems.
-These results hold promise for generating robust and stable entanglement in nonlinear systems, which could prove instrumental for advancements in quantum information processing and quantum computing applications.
+This behavior suggests a process of entanglement generation facilitated by the Duffing nonlinearity.
+Notably, the induced entanglement displayed superior resilience against thermal fluctuations.
+Our findings pave the way for further exploration of entanglement in complex systems.
+These results hold promise for generating strong and stable entanglement in nonlinear systems, which could prove instrumental for advancements in quantum information processing and quantum computing applications.
+Appendix A Discussions on the experimental realization of our proposal
+Our proposal can be realized both in optomechanical and electromechanical platforms.
+For optomechanical structures, our model can be thought as a photonic-crystal optomechanical circuit where both photonic and phononic hopping coupling are intrinsic to the system.
+Such a configuration has been recently experimentally implemented in Fang et al.
+(2017), where a reduced model consists of two optomechanical cavities in which the involved optical (mechanical) resonators are optically (mechanically) coupled through a photonic (phononic) hopping coupling.
+Under a large-detuning regime, one of the optical mode has been adiabatically eliminated, reducing the four mode system to a three-mode loop-coupled optomechanical system consisting of two mechanically coupled mechanical modes which are coupled to a single cavity-field mode as depicted in Figure 1c of our main text.
+Moreover, a driving of the optomechanical cavities with phase-correlated fields results to an effective synthetic magnetic field into the system.
+Similar procedure can be also engineered in a membrane-in-the-middle optomechanical systems, where two mechanically coupled mechanical resonators are located inside an optical cavity as the one shown in Figure 1b of our main text.
+Figure 6: Investigation for nondegenerated mechanical resonators.
+(a) Entanglement between the electromagnetic field and the nonlinear resonator (mode $b_{1}$) versus the synthetic phase $\theta$.
+(b) Entanglement between the electromagnetic field and the linear resonator (mode $b_{2}$) versus the synthetic phase $\theta$.
+The solid line is without Duffing nonlinearity ($\eta=0$), while the dash-dotted and dashed lines are for $\eta=5\times 10^{-6}\omega_{m}$ and $\eta=5\times 10^{-4}\omega_{m}$, respectively.
+We have used $\kappa=2\times 10^{-1}\omega_{m}$, $J_{m}=2\times 10^{-1}\omega_{m}$, $\alpha^{in}=1000\omega_{m}^{1/2}$ and $n_{th}=100$.
+The mechanical frequencies and dissipations are $\omega_{1}=\omega_{m}$, $\gamma_{1}=\gamma_{m}$, $\omega_{2}=\frac{4\omega_{m}}{3}$, and $\gamma_{2}=\frac{4\gamma_{m}}{3}$.
+The other parameters are as in Fig.2.
+Our proposed model can be also implemented in an electromechanical system as the one depicted in Figure 1a of our main text, which consists of multiple mechanical resonators driven by a common microwave cavity represented by an equivalent inductance L and capacitance C.
+The mechanical resonators involved are the moving capacitor gates, which are grounded to the substrate that is seen as the phononic hopping coupling.
+In such a system, the phase-dependent phonon-hopping interaction (i.e., the phase in a loop coupling induces an effective synthetic magnetism) between the nearest-neighboring mechanical resonators can be engineered by coupling them to a superconducting charge qubit.
+More details and calculations on this procedure can be found in Lai et al.
+(2022) of our manuscript.
+It has been assumed in our investigation that the mechanical resonators involved are degenerated, that was to easy both our notations and calculations.
+However, we would like to mention that this assumption does not limit/destroy our results and observed phenomena.
+For that purpose, we have considered nondegenerated mechanical resonators and have displayed in Figure 6 a similar observed behaviors as depicted in Figure 3 of the main text.
+It can be seen that the observed phenomena are the same regardless the physical properties of the involved mechanical resonators.
+Such a conclusion reveals the fact that our investigation is not parameter-limitted, and can be flexibly reproduced experimentally.
 Acknowledgments
@@
 Sarma acknowledges the STARS scheme, MoE, government of India (Proposal ID 2023-0161).
+A.-H.
+Abdel-Aty is thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program.